# Brain Tumor Classification on MRI in Light of Molecular Markers

### v2

@@ Abstract
In research findings, co-deletion of the 1p/19q gene is associated with clinical outcomes in low-grade gliomas. The ability to predict 1p19q status is critical for treatment planning and patient follow-up. This study aims to utilize a specially MRI-based convolutional neural network for brain cancer detection. Although public networks such as RestNet and AlexNet can effectively diagnose brain cancers using transfer learning, the model includes quite a few weights that have nothing to do with medical images. As a result, the diagnostic results are unreliable by the transfer learning model. To deal with the problem of trustworthiness, we create the model from the ground up, rather than depending on a pre-trained model. To enable flexibility, we combined convolution stacking with a dropout and full connect operation, it improved performance by reducing overfitting. During model training, we also supplement the given dataset and inject Gaussian noise. We use three–fold cross-validation to train the best selection model. Comparing InceptionV3, VGG16, and MobileNetV2 fine-tuned with pre-trained models, our model produces better results. On an validation set of 125 codeletion vs. 31 not codeletion images, the proposed network achieves 96.37% percent F1-score, 97.46% percent precision, and 96.34% percent recall when classifying 1p/19q codeletion and not codeletion images.
@@ 1 Introduction
The most efficient method to detect brain tumors is MRI [17, 18]. Scanning generates a vast amount of magnetic resonance images, which is examined by a radiologist.Biomarker detection helps give patients the best appropriate treatment for their particular condition. This study is notable for the uniqueness and promising findings of merging deep learning with radiogenomics. Deep learning was better at detecting 1p/19q co-deletion on T2 images than it does on post-T1 contrast images.
@@
Our contributions are listed below.:
  • Using convolution stack, we develop a CNN specifically for detecting brain cancer in MRI images..
@@ 2 Materials and Methods
To train our proposed network, we use the Kaggle public dataset[10]. Meanwhile, on this dataset, we compare the results of VGG16, InceptionResNetV2, and MobileNetV2 all fine-tuned pre-trained models applying transfer learning.
@@ 3 Results
From Table 4 We discovered that our model is very well balanced regarding Precision, Recall and F1 score, these values are very close. As the comprehensive indicator of Precision and Recall, the F1 score can see the model’s performance in recognition more clearly. In terms of performance,our model exceeds all other models in terms of F1 score.
@@ 4 Discussion
We present a viable CNN-based technique for predicting deletions of the 1p/19q chromosomal arm. In the procedure of applying deep learning algorithms to medical imaging, having a sufficient quantity of datasets is a big challenge. For medical diagnosis, current transfer learning methods leverage a variety of publically available models trained on big public ImageNet datasets. These models, on the other hand, generate a significant quantity of medical images, endangering the accuracy of clinical diagnosis [15].
@@
Small dataset [17] can result in substantial training errors.In order to solve this problem, the training phase, Gaussian noise is included to improve the model’s normalizing ability and fault - tolerant. Adding noise means that the network cannot remember training samples because they are changing all the time, as a result of which network weights are reduced and the network becomes more robust with lower generalization error. Because fresh samples are drawn from the field proximity to existing samples, The input space’s shape has been smoothed out. It make easier for the network to learn the mapping function, resulting in better and faster learning.
@@ 5 Conclusion
In this paper, we use convolution stacking in conjunction with Gaussian noise and random drop in training to create a model for identifying small dataset of brain tumors that outperforms the transfer learning method while having the advantages of small model size and high reliability.Cross-validation in training is an excellent way to overcome the problem of data imbalance.

### v3

@@ Abstract
In research findings, co-deletion of the 1p/19q gene is associated with clinical outcomes in low-grade gliomas. The ability to predict 1p19q status is critical for treatment planning and patient follow-up. This study aims to utilize a specially MRI-based convolutional neural network for brain cancer detection. Although public networks such as RestNet and AlexNet can effectively diagnose brain cancers using transfer learning, the model includes quite a few weights that have nothing to do with medical images. As a result, the diagnostic results are unreliable by the transfer learning model. To deal with the problem of trustworthiness, we create the model from the ground up, rather than depending on a pre-trained model. To enable flexibility, we combined convolution stacking with a dropout and full connect operation, it improved performance by reducing overfitting. During model training, we also supplement the given dataset and inject Gaussian noise. We use three–fold cross-validation to train the best selection model. Comparing InceptionV3, VGG16, and MobileNetV2 fine-tuned with pre-trained models, our model produces better results. On a validation set of 125 codeletion vs. 31 not codeletion images, the proposed network achieves 96.37% percent F1-score, 97.46% percent precision, and 96.34% percent recall when classifying 1p/19q codeletion and not codeletion images.
@@ 1 Introduction
The most efficient method to detect brain tumors is MRI [17, 18, 19]. Scanning generates a vast amount of magnetic resonance images, which is examined by a radiologist. Biomarker detection helps give patients the best appropriate treatment for their particular condition. This study is notable for the unique and promising findings of merging deep learning with radiogenomics. Deep learning was better at detecting 1p/19q co-deletion on T2 images than it does on post-T1 contrast images.
@@
Our contributions are listed below.
  • Using a convolution stack, we develop a CNN specifically for detecting brain cancer in MRI images.
@@ 2 Materials and Methods
To train our proposed network, we use the public Kaggle dataset[10]. Meanwhile, in this data set, we compare the results of VGG16, InceptionResNetV2 and MobileNetV2, all fine-tuned [27] pre-trained models applying transfer learning.
@@ 3 Results
From Table 4, we observe that our model is well-balanced in terms of Precision, Recall, and F1 score, with these values being very close to each other. The F1 score, as a comprehensive metric that combines Precision and Recall, provides a clearer picture of the model’s performance in recognition. In terms of performance, our model surpasses all other models in terms of the F1 score.
@@ 4 Discussion
We present a viable CNN-based technique for predicting deletions of the 1p/19q chromosomal arm. A significant challenge in applying deep learning algorithms to medical imaging is the lack of sufficient datasets. For medical diagnosis, current transfer learning methods rely on various publicly available models trained on large ImageNet datasets. However, these models often generate a substantial number of medical images, which can compromise the accuracy of clinical diagnosis [16].
@@
A small dataset [18] can result in substantial training errors. In order to solve this problem, Gaussian noise is added during the training phase to improve the model’s normalization ability and fault tolerance. Adding noise prevents the network from memorizing the training samples, as they are constantly changing. As a result, the network weights are reduced, making the network more robust and leading to a lower generalization error. Since fresh samples are drawn close to existing samples in the input space, the shape of the input space is smoothed out. This makes it easier for the network to learn the mapping function, resulting in faster and more effective learning. In the next step, we will explore using diffusion [20] models to generate small samples, combine them with LLM for 3D reconstruction [31, 32, 33], and apply compression [21, 22, 23] methods for deployment on embedded devices [24, 25, 26, 29, 30].
@@ 5 Conclusion
In this paper, we use convolution stacking in conjunction with Gaussian noise and random drop in training to create a model for identifying a small dataset of brain tumors that outperforms the transfer learning method while having the advantages of small model size and high reliability. Cross-validation in training is an excellent way to overcome the problem of data imbalance.